# On-device Collaborative Language Modeling via a Mixture of Generalists and Specialists

### v1

@@ Appendix E Visualization of Expert Specialization
To visualize which tokens are routed to the generalist and specialist experts for our 1G1S model trained on SlimPajama, we ask ChatGPT to generate texts in the style of StackExchange, Python Codes, ArXiv Paper and Books. We then feed those texts to the user-specific models and color the token with the Top1 routed index. The routing results after the very first layer (0th), a middle layer (5th), and the very last layer (11th) are presented in Figure 13, 14 and 15.
@@
We perform the same experiments on AG News, asking ChatGPT to generate News text on the topics World, Sports, Business, and Sci/Tech. The routing results after the very first layer (0th), a middle layer (5th), and the very last layer (11th) are presented in Figure 16, 17 and 18.
For all the plots, diagonal entries are in-domain texts and off-diagonal entries are out-of-domain texts.
@@
Figure 13: Visualization of token-level routing results for 1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 0th (first) layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are trained on the categories StackExchange, GitHub, ArXiv, and Book respectively. Texts are generated by ChatGPT.
@@
Figure 14: Visualization of token-level routing results for 1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 5th layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are trained on the categories StackExchange, GitHub, ArXiv, and Book respectively. Texts are generated by ChatGPT.
@@
Figure 15: Visualization of token-level routing results for 1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 11th (last) layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are trained on the categories StackExchange, GitHub, ArXiv, and Book respectively. Texts are generated by ChatGPT.
@@
Figure 16: Visualization of token-level routing results for 1G1S trained on AG News. Tokens are colored with the first expert choice at the 0th (first) layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are with training data from AG News categories World, Sports, Business, and Sci/Tech respectively. Texts are generated by ChatGPT.
@@
Figure 17: Visualization of token-level routing results for 1G1S trained on AG News. Tokens are colored with the first expert choice at the 5th (middle) layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are with training data from AG News categories World, Sports, Business, and Sci/Tech respectively. Texts are generated by ChatGPT.
@@
Figure 18: Visualization of token-level routing results for 1G1S trained on AG News. Tokens are colored with the first expert choice at the 11th (last) layer. Orange denotes the generalist and blue denotes the specialist. Users 0-3 are with training data from AG News categories World, Sports, Business, and Sci/Tech respectively. Texts are generated by ChatGPT.

### v2

@@ Appendix E Visualization of Expert Specialization
To visualize which tokens are routed to the generalist and specialist experts for our CoMiGS-1G1S model trained on SlimPajama, we ask ChatGPT to generate texts in the style of StackExchange, Python Codes, ArXiv Paper and Books. We then feed those texts to the user-specific models and color the token with the Top1 routed index. The routing results after the very first layer (0th), a middle layer (5th), and the very last layer (11th) are presented in Figure 14, 15 and 16.
@@
We perform the same experiments on AG News, asking ChatGPT to generate News text on the topics World, Sports, Business, and Sci/Tech. The routing results after the very first layer (0th), a middle layer (5th), and the very last layer (11th) are presented in Figure 17, 18 and 19.
For all the plots, diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts.
@@
Figure 14: Visualization of token-level routing results for CoMiGS-1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 0th (first) layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.
@@
Figure 15: Visualization of token-level routing results for CoMiGS-1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 5th layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.
@@
Figure 16: Visualization of token-level routing results for CoMiGS-1G1S trained on SlimPajama. Tokens are colored with the first expert choice at the 11th (last) layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.
@@
Figure 17: Visualization of token-level routing results for CoMiGS-1G1S trained on AG News. Tokens are colored with the first expert choice at the 0th (first) layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.
@@
Figure 18: Visualization of token-level routing results for CoMiGS-1G1S trained on AG News. Tokens are colored with the first expert choice at the 5th (middle) layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.
@@
Figure 19: Visualization of token-level routing results for CoMiGS-1G1S trained on AG News. Tokens are colored with the first expert choice at the 11th (last) layer. Orange denotes the generalist and blue denotes the specialist. Diagonal entries are in-distribution texts and off-diagonal entries are out-of-distribution texts. Texts are generated by ChatGPT.